### v1

# Multivariate Fields of Experts

Abstract
We introduce the multivariate fields of experts, a new framework for the learning of image priors. Our model generalizes existing fields of experts methods by incorporating multivariate potential functions constructed via Moreau envelopes of the $\ell_{\infty}$-norm. We demonstrate the effectiveness of our proposal across a range of inverse problems that include image denoising, deblurring, compressed-sensing magnetic-resonance imaging, and computed tomography. The proposed approach outperforms comparable univariate models and achieves performance close to that of deep-learning-based regularizers while being significantly faster, requiring fewer parameters, and being trained on substantially fewer data. In addition, our model retains a relatively high level of interpretability due to its structured design.
Index Terms:
Learnable regularizer, variational methods, inverse problems, Moreau envelope.
11footnotetext: The authors are with the Biomedical Imaging Group, École Polytechnique Fédérale de Lausanne (EPFL), {forename.name}@epfl.ch.
I Introduction
In many areas of science and engineering, it is common to recover an object of interest from indirect linear measurements. Let $\mathbf{H}\in\mathbb{R}^{m\times n}$ denote the measurement operator and let $\mathbf{y}\in\mathbb{R}^{m}$ be the observed data. The goal is to reconstruct the underlying signal $\mathbf{x}\in\mathbb{R}^{n}$ such that $\mathbf{H}\mathbf{x}\approx\mathbf{y}$. However, since $\mathbf{y}$ is usually noisy and $\mathbf{H}$ is often ill-conditioned or even rank-deficient, the direct inversion is typically unstable and results in poor reconstructions. To address this, a popular strategy is to use variational regularization to recover $\mathbf{x}$ as the minimizer of the energy
$\frac{1}{2}\|\mathbf{H}\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x}),$ (1)
where the first term enforces data fidelity and where the regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ encodes prior knowledge about the underlying data.
Many popular regularizers [1, 2, 3, 4, 5, 6, 7] can be expressed as a sum of pointwise potential functions applied to the data convolved with a filter. In particular, the weakly convex ridge regularizer (WCRR) [5] represents a recent advance in this direction, where univariate nonconvex potentials are learned via linear splines.
In this work, we propose the multivariate fields of experts [^1] https://github.com/StanislasDucotterd/MFoE (MFoE), a novel class of learned regularizers that extends the classic fields of experts (FoE) framework by incorporating multivariate potential functions. Our contributions are as follows.
  • We introduce a class of parametric potentials that almost exactly reproduce the spline-based potentials learned in WCRR;
  • we extend WCRR to the multivariate setting by incorporating the $\ell_{\infty}$-norm into this parametric potential framework;
  • we demonstrate the effectiveness of our approach across multiple inverse problems: denoising, deblurring, compressed-sensing magnetic-resonance imaging (CS-MRI), and computed tomography (CT).
II Related Work
A popular model for image reconstruction is the total variation (TV) regularizer [8], where the regularization term $R$ promotes piecewise-constant solutions by penalizing the horizontal and vertical finite differences of the image. Since $R$ is not differentiable, the objective (1) is typically minimized via the proximal operator
$\operatorname{prox}_{\lambda R}(\mathbf{y})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\left(\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x})\right),$ (2)
which yields the solution of a denoising problem—the simplest instance of (1) with $\mathbf{H}=\mathbf{I}$. This proximal operator does not admit a closed-form solution and must be computed using iterative algorithms [9].
To further improve the reconstruction performance of variational methods, many works have proposed data-driven approaches to construct $R$. The FoE framework [1] introduced a regularizer defined, for any image $\mathbf{x}\in\mathbb{R}^{n}$, as
$R(\mathbf{x})=\sum_{k=1}^{K}\langle\mathbf{1}_{n},\psi_{k}(\mathbf{W}_{k}\mathbf{x})\rangle,$ (3)
where $\mathbf{W}_{k}\in\mathbb{R}^{n\times n}$ denotes a convolution matrix and $\psi_{k}\colon\mathbb{R}\to\mathbb{R}$ is a pointwise nonlinearity. In the original FoE model, the potentials $\psi_{k}$ correspond to the negative log-likelihood of Student-$t$ distributions and the filters $\{\mathbf{W}_{k}\}_{k=1}^{K}$ are learned by the minimization of $\sum_{m=1}^{M}R(\mathbf{x}_{m})$ over a dataset $\{\mathbf{x}_{m}\}_{m=1}^{M}$ of natural images .
A more advanced strategy for the learning of the regularizer was proposed in [2], where bilevel optimization [10] is employed. In this setting, the filters are trained to minimize the denoising loss
$\sum_{m=1}^{M}\|\operatorname{prox}_{\lambda R}(\mathbf{x}_{m}+\mathbf{n}_{m})-\mathbf{x}_{m}\|_{2}^{2},$ (4)
where $\mathbf{n}_{m}$ is additive white Gaussian noise. Although the potentials $\{\psi_{k}\}_{k=1}^{K}$ are generally nonconvex in FoE, which makes the proximal operator $\operatorname{prox}_{\lambda R}$ potentially set-valued, bilevel optimization remains applicable and effective in practice.
Subsequent works [3, 4, 5, 6, 7] have further increased expressivity of the FoE model by jointly learning the filters and the nonlinearities $\{\psi_{k}\}_{k=1}^{K}$. While the WCRR learns the gradient of $R$ by parameterizing the derivatives of univariate potentials with linear splines, our model extends the approach to the multivariate setting using higher-dimensional potential functions.
Another important framework for image reconstruction is plug-and-play (PnP) [11], where an off-the-shelf denoiser is inserted as a surrogate for the proximal operator in iterative optimization algorithms. While PnP methods are inspired by variational formulations, they generally do not correspond to the minimization of a well-defined energy functional and therefore lack convergence guarantees in general. Several strategies have been proposed to address this issue, such as Lipschitz constraints on the denoiser [12, 13], or continuation schemes [14] that gradually reduce the influence of the denoiser over the course of optimization.
Finally, a recent line of work have bridged the gap between variational methods and the PnP framework by learning an explicit regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ using deep neural networks [15, 16, 17]. These methods supervise the model via its gradient $\nabla R$, which requires the computation of second-order derivatives during training. This yields highly expressive regularizers with an explicit objective and general convergence guarantees, but comes at the cost of significant computational overhead—both during training, due to the need to differentiate the objective twice, and at inference time, where the gradient of the neural network must be evaluated at each optimization step.
III Description of the Model
III-A Notations and Definitions
The least-squares projection onto a set $\mathcal{C}$ is defined as
$\operatorname{Proj}_{\mathcal{C}}(\mathbf{x})=\operatorname*{arg\,min}_{\mathbf{z}\in\mathcal{C}}\|\mathbf{z}-\mathbf{x}\|_{2}^{2}.$ (5)
This projection exists and is unique whenever the set $\mathcal{C}$ is closed and convex. A particular closed convex set of interest in this work is the $\ell_{1}$-ball in $\mathbb{R}^{d}$, which we denote by
$\mathcal{B}_{\ell_{1}}=\left\{\mathbf{z}\in\mathbb{R}^{d}:\|\mathbf{z}\|_{1}\leq 1\right\}.$ (6)
The Moreau envelope [18] of a proper, lower semi-continuous, convex function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ with parameter $\mu>0$ is defined as
$M_{f}^{\mu}(\mathbf{x})=\inf_{\mathbf{z}\in\mathbb{R}^{d}}\left(f(\mathbf{z})+\frac{1}{2\mu}\|\mathbf{z}-\mathbf{x}\|_{2}^{2}\right).$ (7)
The gradient of the Moreau envelope of $f$ can be related to the proximal operator of $f$ through the equality
$\nabla M_{f}^{\mu}(\mathbf{x})=\frac{1}{\mu}\left(\mathbf{x}-\operatorname{prox}_{\mu f}(\mathbf{x})\right).$ (8)
In this paper, we denote the Moreau envelope (with parameter $\mu$) of the $\ell_{\infty}$-norm in $\mathbb{R}^{d}$ as $\rho_{d}^{\mu}\colon\mathbb{R}^{d}\to\mathbb{R}_{+}$. Using the definition in (7) and the identity
$\operatorname{prox}_{\mu\|\cdot\|_{\infty}}(\mathbf{x})=\mathbf{x}-\mu\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu)$ (9)
with $\mathbf{x}\in\mathbb{R}^{d}$, we can write that
$\rho_{\mu}^{d}(\mathbf{x})=\|\mathbf{x}-\mu\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu)\|_{\infty}+\frac{\mu}{2}\|\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu)\|_{2}^{2}.$ (10)
From (8), we then obtain the gradient
$\nabla\rho_{\mu}^{d}(\mathbf{x})=\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu),$ (11)
which is computed via sorting using the Condat algorithm [19].
III-B Generalization of the WCRR
In Figure 1, we illustrate that the potential learned via linear splines with more than 100 knots in the WCRR can be approximated almost to perfection by a simple parametric form with three degrees of freedom: Indeed, the function $t\mapsto\lambda(\rho_{1}^{\mu}(t)-\rho_{1}^{\nu}(t))$ with $\mu=0.002$ and $\lambda=\nu=0.098$ is visually indistinguishable from the WCRR of Figure 1. This type of potential is close to the minimax-concave penalty introduced in [20]. This observation motivates the use of $\rho_{d}^{\mu}$ as a foundation for a multivariate extension, with the property that the model can reduce to the WCRR when $d=1$.
Note that the absolute value is the only one-dimensional function (up to scaling) that is a norm. This suggests to extend it by the consideration of the Moreau envelopes of general norms. We focus on norms for which both the Moreau envelope and its gradient can be computed efficiently. Among the classic choices, the $\ell_{2}$ and $\ell_{\infty}$ norms are particularly tractable. We opt for the $\ell_{\infty}$ norm for two main reasons: first, it empirically leads to better reconstruction results; second, as established in [21], any norm can be approximated arbitrarily well by the composition of the $\ell_{\infty}$ norm with a linear transformation. This makes it a particularly expressive and computationally attractive choice for multivariate potentials.
The WCRR constrains its potentials to be 1-weakly convex to ensure overall convexity during denoising, given the 1-strong convexity of the data term. While this has minimal impact in the univariate case, it limits performance in the multivariate setting. We therefore relax this constraint in our model and allow potentials with larger weak-convexity constants. To isolate the effect of our multivariate formulation from that of the relaxation, we also compare against a relaxed version of WCRR.
Figure 1: Top: Potential learned in WCRR, vertical lines mark the spline control points. Bottom: Difference between two Moreau envelopes of the absolute value with different parameters.
III-C General Setting
We now move beyond the one-dimensional setting and extend the FoE framework to the multivariate case, where the regularizer is defined, for any image $\mathbf{x}\in\mathbb{R}^{n}$, as
$R(\mathbf{x})=\sum_{k=1}^{K}\left\langle\mathbf{1}_{n},\psi_{k}^{d}\left(\mathbf{W}_{k}^{d}\mathbf{x}\right)\right\rangle,$ (12)
where $\mathbf{W}_{k}^{d}\in\mathbb{R}^{(dn)\times n}$ denotes a convolution matrix that maps a single-channel image to a $d$-channel one, and $\psi_{k}^{d}\colon\mathbb{R}^{d}\to\mathbb{R}$ is a nonlinearity given by
$\psi_{k}(\mathbf{x})=\mu_{k}\rho_{\mu_{k}}^{d}(\mathbf{x})-\mu_{k}\rho_{\tau_{k}\mu_{k}}^{d}(\mathbf{Q}_{k}\mathbf{x}),$ (13)
where $\mathbf{Q}_{k}\in\mathbb{R}^{d\times d}$ and $\tau_{k}$ is a scalar.
Theorem 1
If $\mathbf{Q}_{k}$ is such that $\|\mathbf{Q}_{k}\|_{\infty}\leq 1$ and $\tau_{k}>\|\mathbf{Q}_{k}\|_{2}^{2}$, then the nonlinearity defined in (13) is nonnegative with a unique global minimum at the origin and has nonexpansive gradients.
Theorem 1 guarantees that the regularizer is bounded from below and preserves a similar interpretation to the classic FoE, where weak filter responses are favored while stronger responses are increasingly penalized. The proof can be found in Appendix -A.
The gradient of our regularizer can be written as
$\mathbf{\nabla}R(\mathbf{x})=\mathbf{W}^{T}\varphi(\mathbf{Wx}),$ (14)
where $\mathbf{W}\in\mathbb{R}^{(Kdn)\times n}$ refers to the concatenation of all the $d$-channel convolutions $\mathbf{W}_{k}^{d}$ and $\varphi$ refers to the concatenation of all the gradients $\mathbf{\nabla}\psi_{k}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$. The spectral norm of the filters $\|\mathbf{W}\|_{2}$ is enforced to be one via the same methods as in [5], which ensures that $\mathbf{\nabla}R$ is nonexpansive.
We train the filters $\mathbf{W}$, the $K$ matrices $\mathbf{Q}_{k}$, scalars $\mu_{k},\tau_{k}$, and the regularizing strength $\lambda$ to minimize the denoising loss (4).
To efficiently train filters with a large receptive field, we decompose $\mathbf{W}$ into a composition of three zero-padded convolutions with small kernels and an increasing number of output channels as in [4]. Similarly to [22], the filters are constrained to have zero mean.
III-D Optimization
We minimize the objective (1) using accelerated gradient descent with restart based on the value of the objective as described in [23]. We give the pseudocode for our implementation in Algorithm 1 where $f$ corresponds to (1). Throughout the paper, we set
$\gamma=\frac{1}{\|\mathbf{H}\|_{2}^{2}+\lambda}.$ (15)
Since the regularizer has a gradient $\mathbf{\nabla}R$ that is nonexpansive, the Lipschitz constant of the gradient of (1) is upper-bounded by the Lipschitz constant of the data-fitting term plus the regularization parameter $\lambda$.
The computational cost of the fourth line of Algorithm 1 turns out to be negligible, due to the nature of our potentials. Indeed, from (10) and (11), we have that
$\rho_{d}^{\mu}(\mathbf{x})=\|\mathbf{x}-\mu\nabla\rho_{d}^{\mu}(\mathbf{x})\|_{\infty}+\frac{\mu}{2}\|\nabla\rho_{d}^{\mu}(\mathbf{x})\|_{2}^{2},$ (16)
which makes it very easy to compute the objective on top of the gradient. Therefore, the restart strategy allows us to decrease the amount of iterations without increasing the time required to do one.
Algorithm 1 Accelerated gradient descent with restart based on the objective
1: Input: Function $f$, stepsize $\gamma$, initialization $\mathbf{x}_{0}\in\mathbb{R}^{d}$, tolerance $\varepsilon>0$, number of iterations $n_{\text{max}}$
2:Set $t_{0}=1$, $k=0$, $r_{0}=\infty$, $f_{0}=\infty$, $\mathbf{z}_{0}=\mathbf{x}_{0}$
3:while $r_{k}>\varepsilon$ and $k\leq n_{\text{max}}$ do
4:  $\mathbf{x}_{k+1}=\mathbf{z}_{k}-\gamma\nabla f(\mathbf{z}_{k})$
5:  $f_{k+1}=f(\mathbf{z}_{k})$
6:  $t_{k+1}=\frac{1+\sqrt{1+4t_{k}^{2}}}{2}$
7:  $\mathbf{z}_{k+1}=\mathbf{x}_{k+1}+\frac{t_{k}-1}{t_{k+1}}(\mathbf{x}_{k+1}-\mathbf{x}_{k})$
8:  if $f_{k+1}>f_{k}$ then
9:   $\mathbf{z}_{k+1}=\mathbf{x}_{k+1}$
10:   $t_{k+1}=1$
11:  end if
12:  $r_{k+1}=\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|/\|\mathbf{x}_{k+1}\|$
13:  $k\leftarrow k+1$
14:end while
15: Output: Approximate solution $\mathbf{x}_{k+1}$
III-E Dependence on the Noise Level
To make our regularizer more universal, we train it on a range of Gaussian-noise levels and denote by $R_{\sigma}$ the regularizer associated with the noise level $\sigma$. To keep the model as simple as possible, we share the filters and most other parameters across noise levels, allowing only $\{\mu_{k}\}_{k=1}^{K}$ to depend on $\sigma$. We define a function $\boldsymbol{\mu}\colon\mathbb{R}\to\mathbb{R}^{K}$ that concatenates the $\mu_{k}$, such that $\mu_{k}(\sigma)=\mu(\sigma)_{k}$. This function is parameterized as
$\boldsymbol{\mu}(\sigma)=\frac{1}{100}\left(\sigma\cdot\mathbf{1}_{K}+\frac{1}{20}\mathbf{f}(\bar{\sigma})\right)_{+}+10^{-9}\cdot\mathbf{1}_{K},$ (17)
where $\mathbf{f}\colon\mathbb{R}\to\mathbb{R}^{K}$ is a three-layer neural network. The function $(\cdot)_{+}\coloneqq\max(0,\cdot)$ and the addition of a small term are here to ensure that components of $\boldsymbol{\mu}$ stay positive. The term $\bar{\sigma}$ refers to a min-max normalization of $\sigma$ such that it is in the interval [2,2][-2,2] in order to fully benefit from the neural network $\mathbf{f}$ whose knots at initialization are mostly located in that range.
IV Experiments
IV-A Training on Denoising
We denote by $\boldsymbol{\theta}$ the vector of all parameters in our model. The paired data used for training are $(\mathbf{x}_{m},\mathbf{y}_{m})_{m=1}^{M}$, where each noisy observation is generated as $\mathbf{y}_{m}=\mathbf{x}_{m}+\sigma_{m}\mathbf{n}_{m}$ with Gaussian noise $\mathbf{n}_{m}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and noise levels $\sigma_{m}\sim\mathcal{U}(0,0.2)$. The model is trained to minimize the reconstruction loss
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{M}\sum_{m=1}^{M}\frac{1}{\sqrt{\sigma_{m}}}\|\mathbf{x}_{m}^{*}-\mathbf{x}_{m}\|_{1},$ (18)
where $\mathbf{x}_{m}^{*}\in\operatorname{prox}_{\lambda R_{\sigma_{m}}}(\mathbf{y}_{m})$. The weighting by $1/\sqrt{\sigma_{m}}$ ensures that the model gives sufficient importance to low noise levels during training.
We follow a bilevel strategy to optimize (18). The inner optimization corresponds to the solution $\mathbf{x}^{*}_{m}$ to the proximal problem in (2), as found through accelerated gradient descent with restarts, as described in Algorithm 1. The inner loop terminates when the relative change of the iterates falls below $10^{-4}$ or after 300 iterations.
The outer optimization aims to learn $\boldsymbol{\theta}$ by minimizing the empirical loss (18). This is done using gradient-based optimization with implicit differentiation to compute the required gradients within the inner-optimization process. We use the torchdeq library [24] and apply 25 steps of the Broyden algorithm [25] to approximate the gradient of the fixed-point solution.
The training is conducted on a dataset of 238,400 small patches of size $\left(40\times 40\right)$, extracted from 400 images from the BSD500 dataset [26] and scaled to $[0,1]$. This number of training images is relatively small compared to typical deep-learning-based methods, which often rely on significantly bigger datasets to achieve good generalization [27]. We use the ADAM optimizer with a batch size of 128, setting the learning rate to 0.005 for the filters $\mathbf{W}$ and the function $\boldsymbol{\mu}$, while all other parameters use a learning rate of 0.05. The model is trained for 5000 steps, and the learning rates are decayed by a factor of 0.75 every 500 steps. The full training procedure takes around 5.5 hours on a Tesla V100 GPU.
We report the denoising performance of MFoE in Table I, using $K=15$ potentials of input dimension $d=4$ with filters of size $\left(11\times 11\right)$. For comparison purposes, we include TV [8], the WCRR [5] and Prox-DRUNet [16]. Note that the WCRR results differ slightly from those in the original paper, as we trained it on noise levels ranging from 0 to $50/255$, whereas the original implementation was ranging from 0 to $30/255$. We achieved this by modifying only the splines $s_{\alpha_{i}}$ from [5, Equation (17)], increasing the number of knots from 11 to 18 to maintain interval lengths as close as possible to the original configuration.
As mentioned in Section III-B, we add the relaxed version of WCRR (WCRR-free) and an alternative version of our model based on the $\ell_{2}$-norm (MFoE-$\ell_{2}$). More details about their implementation can be found in Appendices -B and -C.
Relaxing the 1-weak convexity constraint in WCRR does not yield performance improvements, except at the highest noise level. In contrast, the proposed MFoE model consistently outperforms the univariate baselines (WCRR and WCRR-free) as well as the multivariate variant MFoE-$\ell_{2}$. Notably, all four models employ the same number of filters $\left(Kd=60\right)$ and have roughly the same number of parameters, which ensures that the observed performance gains are not due to increased model capacity but result instead from the greater expressiveness of the proposed multivariate potentials.
We present in Table II the performance of the MFoE as we increase the number of filters from $Kd=60$ to 240. Unlike the WCRR, which saturates at 60 filters, our model continues to improve and reaches a performance plateau around 180 filters. To ensure fair comparisons throughout the paper, we adopt the 60-filter variant as our default model.
We further investigate the tradeoff between $K$ and $d$ in Table VII, keeping the total number of filters fixed at $Kd=60$. We observe a significant improvement when moving from $d=1$ to $d=2$, with performance peaking at $d=4$. Beyond that, performance degrades steadily, with $d=60$ performing worse than even the univariate case $d=1$.
TABLE I: Denoising performance in terms of noise level on the BSD68 test set.
$\sigma=15/255$ $\sigma=25/255$ $\sigma=50/255$ Parameters
TV 29.90 27.48 24.83 1
WCRR 31.20 28.68 25.68 $1.4\cdot 10^{4}$
WCRR-free 31.18 28.68 25.78 $1.4\cdot 10^{4}$
MFoE-$\ell_{2}$ 31.22 28.71 25.80 $1.4\cdot 10^{4}$
MFoE 31.32 28.84 25.92 $1.4\cdot 10^{4}$
Prox-DRUNet 31.70 29.18 26.13 $1.7\cdot 10^{7}$
TABLE II: Denoising performance in terms of noise level on the BSD68 test set.
$\sigma=15/255$ $\sigma=25/255$ $\sigma=50/255$
MFoE ($K=15$, $d=4$) 31.32 28.84 25.92
MFoE ($K=30$, $d=4$) 31.36 28.88 25.97
MFoE ($K=45$, $d=4$) 31.38 28.90 25.99
MFoE ($K=60$, $d=4$) 31.39 28.90 25.97
IV-B Visualization and Analysis of the Learned Model
We display the filters of the $(K=15,d=4)$ model in Figure 2. Filters within the same group tend to exhibit similar and complementary structures, while filters across different groups often vary significantly. Since the potential function depends primarily on the filter that produces the strongest response, this behavior suggests a form of invariance in the model—possibly to orientation, scale, or other transformations.
Figure 2: Learned filters on the denoising task with $K=15$ and $d=4$. The black lines split the filters into the 15 different groups of 4.
Figure 3: Denoising results for $\sigma=0.2$, shown on the full image and on a zoomed-in textured region. The reported metrics are (PSNR, SSIM).
We also plot a learned potential of the $(K=30,d=2)$ model along with its two corresponding filter in Figure 4 and show two additional ones in Section -E.
Figure 4: Learned potential function and the corresponding filters.
A direct corollary of Theorem 1 is that the minimizers of our regularizer lie in the null space of $\mathbf{W}$. In Figure 5, we show the impulse response of the filter that corresponds to $\mathbf{W}^{T}\mathbf{W}$, along with the magnitude of its Fourier transform. While the impulse response is close to a Kronecker impulse, it cannot be exactly so as it must sum up to zero due to the zero-mean constraint on all filters in $\mathbf{W}$. The frequency response reveals that only the zero-frequency coefficient vanishes, which suggests that the null space of $\mathbf{W}$ is one-dimensional and consists of constant images. However, this is not entirely accurate: because our filters are zero-padded, $\mathbf{W}^{T}\mathbf{W}$ is not perfectly diagonalized by the discrete Fourier transform—although the approximation is very close, as shown in [28]. It is also intuitive that nonzero constant images are not in the null space of $\mathbf{W}$, as the discontinuities they create at the boundaries result in nonzero filter responses.
To complete the analysis, we run power iterations on $\left(\mathbf{I}-\mathbf{W}^{T}\mathbf{W}\right)$ in float64 precision to accurately estimate the smallest singular value of $\mathbf{W}$. The result, in excess of $2.5\cdot 10^{-3}$, shows that the null space of $\mathbf{W}$ is empty. Consequently, our regularizer admits a unique global minimum, the zero-valued image.
A qualitative comparison of denoising results between our model and univariate baselines reveals that our approach performs particularly well in high-frequency regions that resemble textures as shown in Figure 3. This observation aligns with techniques used in classic texture-classification frameworks [29, 30], where features are derived from the strongest response across a bank of filters with varying orientations and scales.
Figure 5: Impulse and frequency response of $\mathbf{W}^{T}\mathbf{W}$. The image size on the left is $\left(21\times 21\right)$, identical to the size of $\mathbf{W}^{T}\mathbf{W}$. The image size on the right is $\left(1500\times 1500\right)$, which accomodates for some amount of zero-padding in the Fourier transform.
IV-C Inverse Problems
We benchmark MFoE for image deblurring, CS-MRI, and CT. For this, we simulate measurements as
$\mathbf{y}=\mathbf{Hx}+\mathbf{w},$ (19)
where $\mathbf{H}$ is the linear operator that corresponds to the imaging modality and $\mathbf{w}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathbf{w}}\mathbf{I})$.
We compare our learned regularizer with TV, the WCRR, and Prox-DRUNet. Image reconstruction with these methods requires the minimization of the objective (1). This is done via an iterative optimization algorithm which is terminated when the relative difference of the iterates is below $10^{-5}$ or when the number of iterations exceeds 1000.
To deploy the models, we tune their hyperparameters on a validation set using the coarse-to-fine grid search from [4]. Then, the performance is reported on a dedicated test set. For TV, we only tune the regularization strength $\lambda$ in (1). For the other methods, we tune $\lambda$ and the noise level $\sigma$.
IV-C1 Deblurring
We evaluate the frameworks on a deblurring task, where $\mathbf{H}$ is a convolution operator. We took from [31] three kernels of size (25 $\times$ 25) shown on the top of Table III. One kernel is isotropic Gaussian and the other two are anisotropic kernels that correspond to a motion blur. We evaluate the noise levels $\sigma_{\mathbf{w}}\in\{0.01,0.03\}$. Each method is initialized with the observed blurred and noisy measurements, the hyperparameters are tuned on the Set12 dataset, and the performances are reported on the BSD68 dataset.
We report the performances in Table III. The MFoE outperforms the WCRR in every instance and manages to match the performances of Prox-DRUNet for the isotropic blur. Examples of deblurring are provided in Figure 6. There, we can see in the zoomed part that Prox-DRUNet makes a somewhat bolder prediction than the other models. This results in a higher SSIM but a slightly lower PSNR than MFoE, despite its blurrier prediction.
TABLE III: PSNR values for the deblurring on the BSD68 dataset.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$\sigma_{\mathbf{w}}=0.01$ $\sigma_{\mathbf{w}}=0.03$ $\sigma_{\mathbf{w}}=0.01$ $\sigma_{\mathbf{w}}=0.03$ $\sigma_{\mathbf{w}}=0.01$ $\sigma_{\mathbf{w}}=0.03$
TV 26.74 25.62 29.16 26.17 29.36 26.05
WCRR 27.25 26.09 30.26 26.88 30.38 26.75
MFoE 27.36 26.18 30.43 27.14 30.65 27.04
Prox-DRUNet 27.34 26.17 30.55 27.24 30.78 27.15
Figure 6: Reconstruction for a blur with the third kernel and $\sigma_{\mathbf{w}}=0.03$. The reported metrics are (PSNR, SSIM).
IV-C2 Compressed-Sensing MRI
We use fully sampled knee images of size $\left(320\times 320\right)$ from the fastMRI dataset [32] as ground truth, including scans with (PDFS) and without (PD) fat suppression. The corresponding measurements are obtained by a subsampling of the 2D Fourier transform ($k$-space) using a Cartesian sampling mask defined by two parameters: the acceleration factor $M_{\text{acc}}$ and the center fraction $M_{\text{cf}}$. Specifically, the central $\lfloor 320M_{\text{cf}}\rfloor$ columns of $k$-space (those corresponding to low frequencies) are fully retained, while the remaining columns are uniformly subsampled, which results in a total of $\lfloor 320/M_{\text{acc}}\rfloor$ selected columns. Gaussian noise with standard deviation $\sigma_{\mathbf{w}}=0.01$ is added to the real and the imaginary part of the measurements. Each method is initialized with the zero-filled image, the hyperparameters are tuned on a validation set of 10 images, and the performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
The performance is reported in Table IV. Again, the MFoE outperforms the WCRR in every instance. It also manages to outperform Prox-DRUNet on the simpler instance with the PD data. Examples of reconstruction are provided in Figure 7.
TABLE IV: PSNR values for MRI reconstruction on the fastMRI dataset.
$\left(M_{\text{acc}}=4,M_{\text{cf}}=0.08\right)$ $\left(M_{\text{acc}}=8,M_{\text{cf}}=0.04\right)$
PD PDFS PD PDFS
Zero-fill 27.43 29.62 23.49 27.00
TV 33.23 32.70 26.97 29.42
WCRR 35.10 34.18 29.32 31.00
MFoE 35.40 34.21 30.26 31.20
Prox-DRUNet 35.32 34.53 30.97 31.66
Figure 7: Reconstruction for 8-fold subsampling and PDFS data. The reported metrics are (PSNR, SSIM).
IV-C3 Computed Tomography
We evaluate our method on the LoDoPaB-CT dataset [33], which contains ground-truth images derived from 2D slices of reconstructions in the LIDC/IDRI database [34], each of size $\left(362\times 362\right)$. The measurement operators are defined by parallel-beam geometry and implemented using the ASTRA Toolbox [35, 36]. All experiments are conducted with 256 simulated detectors. The resulting sinograms are corrupted with Gaussian noise of standard deviation $\sigma_{\mathbf{w}}=0.1$.
We consider three acquisition settings with 60, 40, and 20 projection angles, respectively. Each method is initialized with the filtered-back projected (FBP) image, and the hyperparameters are tuned on a validation set of 10 images. The performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
The performance is presented in Table V. There too, MFoE outperforms WCRR in all tested configurations. Examples of reconstruction are provided in Figure 8.
TABLE V: PSNR values for CT reconstruction on the LoDoBaP dataset.
60 angles 40 angles 20 angles
FBP 28.53 25.87 21.41
TV 33.74 32.58 30.11
WCRR 35.21 33.87 31.08
MFoE 35.45 34.22 31.54
Prox-DRUNet 35.73 34.61 32.12
Figure 8: Reconstruction for CT with 20 angles and 256 line detectors. The reported metrics are (PSNR, SSIM).
IV-C4 Time Comparisons
Table VI contains the average inference time per image, measured in seconds, for each inverse problem. More detailed results are provided in Tables $X$. These results highlight that the performance gains of Prox-DRUNet come at a computational cost. In particular, the CT reconstructions with Prox-DRUNet are significantly slower than with other methods, as they require the repeated evaluation of the proximal operator of the data-fitting term—a step that is especially costly in the case of CT. By contrast, MFoE is only moderately more demanding than the WCRR in inference time, between 35% and 40% on average across all experiments due to the computational cost of the projection onto $\mathcal{B}_{\ell_{1}}$.
TABLE VI: Average time per image for all the reconstruction modalities.
Deblurring CS-MRI CT
TV 5.86 7.16 24.26
WCRR 6.45 6.89 13.90
MFoE 5.39 16.48 16.60
Prox-DRUNet 39.82 62.52 267.42
Deep-learning-based methods such as Prox-DRUNet sets a high standard in learned regularization by leveraging a large number of parameters and extensive training data which contribute to their good performance across diverse tasks. We believe that MFoE offers an interesting alternative with significantly fewer parameters, fewer training data, faster inference time, and with a reconstruction quality that is close to Prox-DRUNet. Together, these methods offer complementary strengths, providing valuable options suited to varying application needs.
V Conclusion
We have proposed the MFoE model. It is a principled extension of the classic FoE model and incorporates multivariate potential functions via Moreau envelopes of the $\ell_{\infty}$-norm. Our method offers improved adaptability across multiple inverse problems. The image quality that it achieves is comparable to that of state-of-the-art learned regularizers. It offers a clear structured design that is easy to implement with competitive reconstruction times.
Acknowledgements
The research leading to these results was supported by the European Research Council (ERC) under European Union’s Horizon 2020 (H2020), Grant Agreement - Project No 101020573 FunLearn.

### v2

@@
-Multivariate Fields of Experts
+Multivariate Fields of Experts for Convergent Image Reconstruction
 Abstract
@@
 The proposed approach outperforms comparable univariate models and achieves performance close to that of deep-learning-based regularizers while being significantly faster, requiring fewer parameters, and being trained on substantially fewer data.
-In addition, our model retains a relatively high level of interpretability due to its structured design.
-Index Terms:
-Learnable regularizer, variational methods, inverse problems, Moreau envelope.
+In addition, our model retains a high level of interpretability due to its structured design.
+It is supported by theoretical convergence guarantees which ensure reliability in sensitive reconstruction tasks.
 11footnotetext: The authors are with the Biomedical Imaging Group, École polytechnique fédérale de Lausanne (EPFL), {forename.name}@epfl.ch.
 I Introduction
-In many areas of science and engineering, it is common to recover an object of interest from indirect linear measurements.
+A common problem in science and engineering is to recover an object of interest from indirect linear measurements.
 Let $\mathbf{H}\in\mathbb{R}^{m\times n}$ denote the measurement operator and let $\mathbf{y}\in\mathbb{R}^{m}$ be the observed data.
@@
 To address this, a popular strategy is to use variational regularization to recover $\mathbf{x}$ as the minimizer of the energy
-$\frac{1}{2}\|\mathbf{H}\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x}),$ (1)
-where the first term enforces data fidelity and where the regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ encodes prior knowledge about the underlying data.
-Many popular regularizers [1, 2, 3, 4, 5, 6, 7] can be expressed as a sum of pointwise potential functions applied to the data convolved with a filter.
-In particular, the weakly convex ridge regularizer (WCRR) [5] represents a recent advance in this direction, where univariate nonconvex potentials are learned via linear splines.
-In this work, we propose the multivariate fields of experts [^1] https://github.com/StanislasDucotterd/MFoE (MFoE), a novel class of learned regularizers that extends the classic fields of experts (FoE) framework by incorporating multivariate potential functions.
+$f(\mathbf{x})=\tfrac{1}{2}\|\mathbf{H}\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x}),$ (1)
+where the first term enforces data fidelity and the regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ encodes prior knowledge about the underlying signal.
+Many popular regularizers [34, 10, 27, 18, 19, 45, 32] can be expressed as a sum of univariate potential functions applied to filter responses, a structure frequently referred to as the fields of experts (FoE) model.
+In particular, the weakly convex ridge regularizer (WCRR) [19] represents a recent advance in this direction, where nonconvex potentials are learned via linear splines.
+While highly effective, such approaches implicitly assume independence between the filter responses (channels), ignoring potentially valuable interactions between them.
+In this work, we propose the multivariate fields of experts (MFoE), a novel class of learned regularizers that generalizes the classic FoE framework by incorporating multivariate potential functions that have the capability to capture interactions between channels.
 Our contributions are as follows.
-  • We introduce a class of parametric potentials that almost exactly reproduce the spline-based potentials learned in WCRR;
-  • we extend WCRR to the multivariate setting by incorporating the $\ell_{\infty}$-norm into this parametric potential framework;
-  • we demonstrate the effectiveness of our approach across multiple inverse problems: denoising, deblurring, compressed-sensing magnetic-resonance imaging (CS-MRI), and computed tomography (CT).
+  • Multivariate Generalization: We extend the WCRR framework to the multivariate setting by introducing a novel class of parametric potentials based on Moreau envelopes.
+  • Tailored Optimization: We design a specialized optimization algorithm for the resulting objective and prove its convergence to a stationary point.
+  • Comprehensive Validation: We demonstrate the effectiveness of our approach across diverse inverse problems that include denoising, deblurring, compressed-sensing magnetic-resonance imaging (CS-MRI), and computed tomography (CT).
+We demonstrate superior performance to univariate FoE methods.
+Our code is accessible on Github  [^1] https://github.com/StanislasDucotterd/MFoE.
 II Related Work
-A popular model for image reconstruction is the total variation (TV) regularizer [8], where the regularization term $R$ promotes piecewise-constant solutions by penalizing the horizontal and vertical finite differences of the image.
-Since $R$ is not differentiable, the objective (1) is typically minimized via the proximal operator
-$\operatorname{prox}_{\lambda R}(\mathbf{y})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\left(\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R(\mathbf{x})\right),$ (2)
+A popular model for image reconstruction is the total variation (TV) regularizer [35], where the regularization term $R_{\text{TV}}$ promotes piecewise-constant solutions by penalizing the horizontal and vertical finite differences of the image.
+Since $R_{\text{TV}}$ is not differentiable, the objective (1) is typically minimized via the proximal operator
+$\operatorname{prox}_{\lambda R}(\mathbf{y})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\left(\tfrac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda R_{\text{TV}}(\mathbf{x})\right),$ (2)
 which yields the solution of a denoising problem—the simplest instance of (1) with $\mathbf{H}=\mathbf{I}$.
-This proximal operator does not admit a closed-form solution and must be computed using iterative algorithms [9].
-To further improve the reconstruction performance of variational methods, many works have proposed data-driven approaches to construct $R$.
-The FoE framework [1] introduced a regularizer defined, for any image $\mathbf{x}\in\mathbb{R}^{n}$, as
-$R(\mathbf{x})=\sum_{k=1}^{K}\langle\mathbf{1}_{n},\psi_{k}(\mathbf{W}_{k}\mathbf{x})\rangle,$ (3)
-where $\mathbf{W}_{k}\in\mathbb{R}^{n\times n}$ denotes a convolution matrix and $\psi_{k}\colon\mathbb{R}\to\mathbb{R}$ is a pointwise nonlinearity.
-In the original FoE model, the potentials $\psi_{k}$ correspond to the negative log-likelihood of Student-$t$ distributions and the filters $\{\mathbf{W}_{k}\}_{k=1}^{K}$ are learned by the minimization of $\sum_{m=1}^{M}R(\mathbf{x}_{m})$ over a dataset $\{\mathbf{x}_{m}\}_{m=1}^{M}$ of natural images .
-A more advanced strategy for the learning of the regularizer was proposed in [2], where bilevel optimization [10] is employed. In this setting, the filters are trained to minimize the denoising loss
-$\sum_{m=1}^{M}\|\operatorname{prox}_{\lambda R}(\mathbf{x}_{m}+\mathbf{n}_{m})-\mathbf{x}_{m}\|_{2}^{2},$ (4)
-where $\mathbf{n}_{m}$ is additive white Gaussian noise.
+This proximal operator does not admit a closed-form solution and must be computed through iterative algorithms [8].
+To further improve the reconstruction performance of variational methods, many works propose data-driven approaches to construct $R$.
+The FoE framework [34] relies on a regularizer defined for any image $\mathbf{x}\in\mathbb{R}^{n}$ as
+$R_{\text{FoE}}(\mathbf{x})=\sum_{k=1}^{K}\langle\mathbf{1}_{n},\psi_{k}(\mathbf{W}_{k}\mathbf{x})\rangle,$ (3)
+where $\mathbf{W}_{k}\in\mathbb{R}^{n\times n}$ denotes a convolution matrix and $\psi_{k}\colon\mathbb{R}\to\mathbb{R}$ is a univariate nonlinearity.
+In the original FoE model, the potentials $\psi_{k}$ correspond to the negative log-likelihood of Student-$t$ distributions and the filters $\{\mathbf{W}_{k}\}_{k=1}^{K}$ are learned by the minimization of $\sum_{m=1}^{M}R_{\text{FoE}}(\mathbf{x}_{m})$ over a dataset $\{\mathbf{x}_{m}\}_{m=1}^{M}$ of natural images.
+A more advanced strategy for the learning of the regularizer is proposed in [10], where bilevel optimization [39] is employed.
+In this setting, the filters are trained explicitly to minimize the denoising loss
+$\sum_{m=1}^{M}\|\operatorname{prox}_{\lambda R_{\text{FoE}}}(\mathbf{x}_{m}+\mathbf{n}_{m})-\mathbf{x}_{m}\|_{2}^{2},$ (4)
+where $\mathbf{n}_{m}$ is an additive white Gaussian noise.
+This method aligns the training objective with the evaluation metric (PSNR), which yields better results than the likelihood-based learning of the original FoE.
 Although the potentials $\{\psi_{k}\}_{k=1}^{K}$ are generally nonconvex in FoE, which makes the proximal operator $\operatorname{prox}_{\lambda R}$ potentially set-valued, bilevel optimization remains applicable and effective in practice.
-Subsequent works [3, 4, 5, 6, 7] have further increased expressivity of the FoE model by jointly learning the filters and the nonlinearities $\{\psi_{k}\}_{k=1}^{K}$.
-While the WCRR learns the gradient of $R$ by parameterizing the derivatives of univariate potentials with linear splines, our model extends the approach to the multivariate setting using higher-dimensional potential functions.
-Another important framework for image reconstruction is plug-and-play (PnP) [11], where an off-the-shelf denoiser is inserted as a surrogate for the proximal operator in iterative optimization algorithms.
-While PnP methods are inspired by variational formulations, they generally do not correspond to the minimization of a well-defined energy functional and therefore lack convergence guarantees in general.
-Several strategies have been proposed to address this issue, such as Lipschitz constraints on the denoiser [12, 13], or continuation schemes [14] that gradually reduce the influence of the denoiser over the course of optimization.
-Finally, a recent line of work have bridged the gap between variational methods and the PnP framework by learning an explicit regularizer $R\colon\mathbb{R}^{n}\to\mathbb{R}$ using deep neural networks [15, 16, 17].
-These methods supervise the model via its gradient $\nabla R$, which requires the computation of second-order derivatives during training.
-This yields highly expressive regularizers with an explicit objective and general convergence guarantees, but comes at the cost of significant computational overhead—both during training, due to the need to differentiate the objective twice, and at inference time, where the gradient of the neural network must be evaluated at each optimization step.
+Subsequent works [27, 18, 19, 45, 32] increase the expressivity of the FoE model further by jointly learning the filters and the nonlinearities $\{\psi_{k}\}_{k=1}^{K}$.
+The WCRR learns the gradient of the regularizer by parameterizing the derivatives of univariate potentials with linear splines.
+Our model relies on higher-dimensional potential functions to extend this approach to the multivariate setting.
+While the approaches discussed above minimize an explicit energy functional, a parallel line of research known as Plug-and-Play (PnP) [43] bypasses the use of an explicit regularizer.
+Instead, PnP methods insert an off-the-shelf denoiser as a surrogate for the proximal operator in iterative optimization algorithms.
+Although PnP yields good performance, it generally does not correspond to the minimization of an explicit objective and often lacks convergence guarantees.
+Several strategies have been proposed to address this issue, such as the enforcement of Lipschitz constraints on the denoiser [36, 13] or the use of continuation schemes [9].
+Finally, a recent line of work has bridged the gap between variational methods and the PnP framework by parameterizing an explicit regularizer with deep neural networks [20, 21, 15].
+These methods learn the regularizer via its gradient, which requires the computation of second-order derivatives during training.
+This yields highly expressive regularizers with an explicit objective and general convergence guarantees, but comes at the cost of a significant computational overhead—both during training, due to the need to differentiate the objective twice, and at inference time, since the gradient of the neural network must be evaluated at each optimization step.
 III Description of the Model
@@
 $\mathcal{B}_{\ell_{1}}=\left\{\mathbf{z}\in\mathbb{R}^{d}:\|\mathbf{z}\|_{1}\leq 1\right\}.$ (6)
-The Moreau envelope [18] of a proper, lower semi-continuous, convex function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ with parameter $\mu>0$ is defined as
+The Moreau envelope [26] of a proper, lower semicontinuous, convex function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ with parameter $\mu>0$ is defined as
 $M_{f}^{\mu}(\mathbf{x})=\inf_{\mathbf{z}\in\mathbb{R}^{d}}\left(f(\mathbf{z})+\frac{1}{2\mu}\|\mathbf{z}-\mathbf{x}\|_{2}^{2}\right).$ (7)
-The gradient of the Moreau envelope of $f$ can be related to the proximal operator of $f$ through the equality
+The gradient of the Moreau envelope of $f$ is related to the proximal operator of $f$ through the equality
 $\nabla M_{f}^{\mu}(\mathbf{x})=\frac{1}{\mu}\left(\mathbf{x}-\operatorname{prox}_{\mu f}(\mathbf{x})\right).$ (8)
-In this paper, we denote the Moreau envelope (with parameter $\mu$) of the $\ell_{\infty}$-norm in $\mathbb{R}^{d}$ as $\rho_{d}^{\mu}\colon\mathbb{R}^{d}\to\mathbb{R}_{+}$.
-Using the definition in (7) and the identity
-$\operatorname{prox}_{\mu\|\cdot\|_{\infty}}(\mathbf{x})=\mathbf{x}-\mu\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu)$ (9)
-with $\mathbf{x}\in\mathbb{R}^{d}$, we can write that
+In this paper, we denote the Moreau envelope (with parameter $\mu$) of the $\ell_{\infty}$-norm in $\mathbb{R}^{d}$ as $\rho^{d}_{\mu}\colon\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$.
+We rely on the standard identity [30, Equation 6.8]
+$\operatorname{prox}_{\mu\|\cdot\|_{\infty}}(\mathbf{x})=\mathbf{x}-\mu\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu).$ (9)
+The proximal operator explicitly realizes the infimum in the definition of the Moreau envelope.
+Therefore, the evaluation of (7) at $\mathbf{z}=\operatorname{prox}_{\mu\|\cdot\|_{\infty}}(\mathbf{x})$ yields
 $\rho_{\mu}^{d}(\mathbf{x})=\|\mathbf{x}-\mu\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu)\|_{\infty}+\frac{\mu}{2}\|\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu)\|_{2}^{2}.$ (10)
-From (8), we then obtain the gradient
+From (8), we then obtain the convenient formula for the gradient
 $\nabla\rho_{\mu}^{d}(\mathbf{x})=\operatorname{Proj}_{\mathcal{B}_{\ell_{1}}}(\mathbf{x}/\mu),$ (11)
-which is computed via sorting using the Condat algorithm [19].
+which can be computed efficiently via sorting using the Condat algorithm [11].
 III-B Generalization of the WCRR
-In Figure 1, we illustrate that the potential learned via linear splines with more than 100 knots in the WCRR can be approximated almost to perfection by a simple parametric form with three degrees of freedom: Indeed, the function $t\mapsto\lambda(\rho_{1}^{\mu}(t)-\rho_{1}^{\nu}(t))$ with $\mu=0.002$ and $\lambda=\nu=0.098$ is visually indistinguishable from the WCRR of Figure 1.
-This type of potential is close to the minimax-concave penalty introduced in [20].
-This observation motivates the use of $\rho_{d}^{\mu}$ as a foundation for a multivariate extension, with the property that the model can reduce to the WCRR when $d=1$.
+The WCRR [19] constructs a regularization functional by applying 1-weakly convex, univariate potential functions parameterized as linear splines to learned filter responses.
+Our proposed model generalizes this framework by replacing these scalar functions with higher-dimensional analogs, a generalization directly motivated by the geometric properties of the WCRR potentials.
+Specifically, these potentials take the form $\psi_{k}(x)=\alpha_{k}^{-2}\psi(\alpha_{k}x)$ where a single base function $\psi$ is shared across all channels.
+In Figure 1, we illustrate that $\psi$ (which is learned via a linear spline with more than 100 knots) can be approximated almost to perfection by a simple three-parameter function of the form $t\mapsto\lambda(\rho_{\mu}^{1}(t)-\rho_{\nu}^{1}(t))$ with $\mu=0.002$ and $\lambda=\nu=0.098$.
+This observation motivates the use of $\rho^{d}_{\mu}$ as a foundation for a multivariate extension, with the property that the model can reduce to the WCRR when $d=1$.
+This three-parameter function is closely related to the minimax-concave penalty introduced in [37].
 Note that the absolute value is the only one-dimensional function (up to scaling) that is a norm.
@@
 We focus on norms for which both the Moreau envelope and its gradient can be computed efficiently.
-Among the classic choices, the $\ell_{2}$ and $\ell_{\infty}$ norms are particularly tractable.
-We opt for the $\ell_{\infty}$ norm for two main reasons: first, it empirically leads to better reconstruction results; second, as established in [21], any norm can be approximated arbitrarily well by the composition of the $\ell_{\infty}$ norm with a linear transformation.
-This makes it a particularly expressive and computationally attractive choice for multivariate potentials.
+The $\ell_{1}$-norm decomposes into a sum of absolute values, which implies that the difference of two Moreau envelopes recovers the univariate formulation of the WCRR.
+The $\ell_{2}$ and $\ell_{\infty}$ norms are particularly tractable.
+We select the $\ell_{\infty}$-norm for two reasons: first, it yields better empirical results; second, the authors of [40, Theorem 4] establish that any norm on $\mathbb{R}^{d}$ can be approximated to arbitrary precision by the function $\mathbf{x}\mapsto\|\mathbf{L}\mathbf{x}\|_{\infty}$, where $\mathbf{L}$ maps into a potentially higher-dimensional space.
+This universality property ensures that our model can represent a broad class of convex potentials simply by learning the appropriate linear filters.
 The WCRR constrains its potentials to be 1-weakly convex to ensure overall convexity during denoising, given the 1-strong convexity of the data term.
-While this has minimal impact in the univariate case, it limits performance in the multivariate setting.
-We therefore relax this constraint in our model and allow potentials with larger weak-convexity constants.
-To isolate the effect of our multivariate formulation from that of the relaxation, we also compare against a relaxed version of WCRR.
-Figure 1: Top: Potential learned in WCRR, vertical lines mark the spline control points. Bottom: Difference between two Moreau envelopes of the absolute value with different parameters.
+Our optimization algorithm allows us to relax this constraint in our model without losing any convergence guarantees.
+Figure 1: The WCRR potential (top) and its gradient (bottom) match the difference of two Moreau envelopes of the absolute value. Vertical ticks indicate the spline knots.
 III-C General Setting
-We now move beyond the one-dimensional setting and extend the FoE framework to the multivariate case, where the regularizer is defined, for any image $\mathbf{x}\in\mathbb{R}^{n}$, as
+We now move beyond the univariate setting and extend the FoE framework to the multivariate case, where the regularizer is defined for any image $\mathbf{x}\in\mathbb{R}^{n}$ as
 $R(\mathbf{x})=\sum_{k=1}^{K}\left\langle\mathbf{1}_{n},\psi_{k}^{d}\left(\mathbf{W}_{k}^{d}\mathbf{x}\right)\right\rangle,$ (12)
-where $\mathbf{W}_{k}^{d}\in\mathbb{R}^{(dn)\times n}$ denotes a convolution matrix that maps a single-channel image to a $d$-channel one, and $\psi_{k}^{d}\colon\mathbb{R}^{d}\to\mathbb{R}$ is a nonlinearity given by
-$\psi_{k}(\mathbf{x})=\mu_{k}\rho_{\mu_{k}}^{d}(\mathbf{x})-\mu_{k}\rho_{\tau_{k}\mu_{k}}^{d}(\mathbf{Q}_{k}\mathbf{x}),$ (13)
+where $\mathbf{W}_{k}^{d}\in\mathbb{R}^{(dn)\times n}$ denotes a multi-convolution matrix that maps a single-channel image to a $d$-channel one, and $\psi_{k}^{d}\colon\mathbb{R}^{d}\to\mathbb{R}$ is a nonlinearity given by
+$\psi_{k}^{d}(\mathbf{x})=\mu_{k}\rho_{\mu_{k}}^{d}(\mathbf{x})-\mu_{k}\rho_{\tau_{k}\mu_{k}}^{d}(\mathbf{Q}_{k}\mathbf{x}),$ (13)
 where $\mathbf{Q}_{k}\in\mathbb{R}^{d\times d}$ and $\tau_{k}$ is a scalar.
 Theorem 1
-If $\mathbf{Q}_{k}$ is such that $\|\mathbf{Q}_{k}\|_{\infty}\leq 1$ and $\tau_{k}>\|\mathbf{Q}_{k}\|_{2}^{2}$, then the nonlinearity defined in (13) is nonnegative with a unique global minimum at the origin and has nonexpansive gradients.
-Theorem 1 guarantees that the regularizer is bounded from below and preserves a similar interpretation to the classic FoE, where weak filter responses are favored while stronger responses are increasingly penalized.
+If $\mathbf{Q}_{k}$ is such that $\|\mathbf{Q}_{k}\|_{\infty}\leq 1$ and $\tau_{k}>\|\mathbf{Q}_{k}\|_{2}^{2}$, then the nonlinearity (13) is nonnegative with a unique global minimum at the origin, and a gradient that is nonexpansive.
+Theorem 1 guarantees that the regularizer is bounded from below and preserves a interpretation similar to the classic FoE, where weak filter responses are favored while stronger responses are penalized.
 The proof can be found in Appendix -A.
@@
 $\mathbf{\nabla}R(\mathbf{x})=\mathbf{W}^{T}\varphi(\mathbf{Wx}),$ (14)
-where $\mathbf{W}\in\mathbb{R}^{(Kdn)\times n}$ refers to the concatenation of all the $d$-channel convolutions $\mathbf{W}_{k}^{d}$ and $\varphi$ refers to the concatenation of all the gradients $\mathbf{\nabla}\psi_{k}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$.
-The spectral norm of the filters $\|\mathbf{W}\|_{2}$ is enforced to be one via the same methods as in [5], which ensures that $\mathbf{\nabla}R$ is nonexpansive.
-We train the filters $\mathbf{W}$, the $K$ matrices $\mathbf{Q}_{k}$, scalars $\mu_{k},\tau_{k}$, and the regularizing strength $\lambda$ to minimize the denoising loss (4).
-To efficiently train filters with a large receptive field, we decompose $\mathbf{W}$ into a composition of three zero-padded convolutions with small kernels and an increasing number of output channels as in [4].
+where $\mathbf{W}\in\mathbb{R}^{(Kdn)\times n}$ refers to the concatenation of all the $d$-channel convolutions and $\varphi$ refers to the concatenation of all the gradients $\mathbf{\nabla}\psi_{k}^{d}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$.
+The spectral norm $\|\mathbf{W}\|_{2}$ of the filters is enforced to be one which ensures that $\mathbf{\nabla}R$ is nonexpansive.
+To efficiently train filters with a large receptive field, we decompose $\mathbf{W}$ into a three zero-padded convolutions with small kernels and an increasing number of output channels, as in [18].
 Similarly to [22], the filters are constrained to have zero mean.
 III-D Optimization
-We minimize the objective (1) using accelerated gradient descent with restart based on the value of the objective as described in [23].
-We give the pseudocode for our implementation in Algorithm 1 where $f$ corresponds to (1).
-Throughout the paper, we set
-$\gamma=\frac{1}{\|\mathbf{H}\|_{2}^{2}+\lambda}.$ (15)
-Since the regularizer has a gradient $\mathbf{\nabla}R$ that is nonexpansive, the Lipschitz constant of the gradient of (1) is upper-bounded by the Lipschitz constant of the data-fitting term plus the regularization parameter $\lambda$.
-The computational cost of the fourth line of Algorithm 1 turns out to be negligible, due to the nature of our potentials. Indeed, from (10) and (11), we have that
-$\rho_{d}^{\mu}(\mathbf{x})=\|\mathbf{x}-\mu\nabla\rho_{d}^{\mu}(\mathbf{x})\|_{\infty}+\frac{\mu}{2}\|\nabla\rho_{d}^{\mu}(\mathbf{x})\|_{2}^{2},$ (16)
-which makes it very easy to compute the objective on top of the gradient. Therefore, the restart strategy allows us to decrease the amount of iterations without increasing the time required to do one.
-Algorithm 1 Accelerated gradient descent with restart based on the objective
-1: Input: Function $f$, stepsize $\gamma$, initialization $\mathbf{x}_{0}\in\mathbb{R}^{d}$, tolerance $\varepsilon>0$, number of iterations $n_{\text{max}}$
-2:Set $t_{0}=1$, $k=0$, $r_{0}=\infty$, $f_{0}=\infty$, $\mathbf{z}_{0}=\mathbf{x}_{0}$
-3:while $r_{k}>\varepsilon$ and $k\leq n_{\text{max}}$ do
-4:  $\mathbf{x}_{k+1}=\mathbf{z}_{k}-\gamma\nabla f(\mathbf{z}_{k})$
-5:  $f_{k+1}=f(\mathbf{z}_{k})$
-6:  $t_{k+1}=\frac{1+\sqrt{1+4t_{k}^{2}}}{2}$
-7:  $\mathbf{z}_{k+1}=\mathbf{x}_{k+1}+\frac{t_{k}-1}{t_{k+1}}(\mathbf{x}_{k+1}-\mathbf{x}_{k})$
-8:  if $f_{k+1}>f_{k}$ then
-9:   $\mathbf{z}_{k+1}=\mathbf{x}_{k+1}$
-10:   $t_{k+1}=1$
-11:  end if
-12:  $r_{k+1}=\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|/\|\mathbf{x}_{k+1}\|$
-13:  $k\leftarrow k+1$
-14:end while
-15: Output: Approximate solution $\mathbf{x}_{k+1}$
-III-E Dependence on the Noise Level
+We design an algorithm that minimizes the objective function defined in (1) and guarantees convergence to a stationary point. We build upon the work of [28] which uses the heavy-ball method [31]
+$\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha_{k}\nabla f(\mathbf{x}_{k})+\beta_{k}(\mathbf{x}_{k}-\mathbf{x}_{k-1}).$ (15)
+The method of [28] ensures convergence by imposing strict constraints on the stepsize $\alpha_{k}$ and momentum $\beta_{k}$.
+Here, we relax these constraints to allow for more aggressive acceleration while still maintaining convergence guarantees, with the backing of Theorem 2 below.
+As described in Algorithm 1 (Lines 5–6), our refinement is a backtracking mechanism that discards the inertial steps and falls back to a standard gradient descent update if the candidate update $\mathbf{x}_{k+1}$ fails to satisfy a sufficient descent condition.
+The backtracking step introduces negligible overhead because the cost of the check of the descent condition is minimal. This is due to the specific structure of our potentials
+$\rho^{d}_{\mu}(\mathbf{x})=\|\mathbf{x}-\mu\nabla\rho^{d}_{\mu}(\mathbf{x})\|_{\infty}+\frac{\mu}{2}\|\nabla\rho^{d}_{\mu}(\mathbf{x})\|_{2}^{2}.$ (16)
+Since $R$ is composed of these potentials, the function value $R(\mathbf{x})$ can be computed cheaply alongside the gradient $\nabla R(\mathbf{x})$.
+Furthermore, the values computed during a rejected candidate step are reused for the fallback step, which ensures that the objective and gradient are evaluated twice only when a restart occurs.
+The nonexpansiveness of $\mathbf{\nabla}R$ implies that $\mathbf{\nabla}f$ is $L$-Lipschitz continuous with constant
+$L=\|\mathbf{H}\|_{2}^{2}+\lambda.$ (17)
+Theorem 2
+Let $\alpha\in(0,2/L)$. If the sequence $\{\mathbf{x}_{k}\}_{k\geq 0}$ generated by Algorithm 1 is bounded, then it converges to a stationary point of $f$ and has a finite length, with
+$\sum_{k\geq 0}\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|_{2}<\infty.$ (18)
+Theorem 2 is essential to the reliability of image reconstruction, as the finite-length property guarantees that the iterates settle to a specific solution rather than being indefinitely oscillating or diverging.
+The proof can be found in Appendix -B.
+The theorem assumes bounded iterates, a condition that can also be satisfied formally by augmenting the objective with a coercive barrier term such as $B(\mathbf{x})=\log\left(1+\exp\left(\|\mathbf{x}\|_{2}^{2}-C\right)\right)$.
+There, the constant term $C$ can be chosen arbitrarily large.
+The coercive term ensures bounded level sets while it remains negligible within the valid range of image intensities.
+Similar ideas have been proposed in [19] and [21].
+Throughout this work, we use the fixed stepsize $\alpha=1.99/L$.
+We increase the momentum parameter $\beta_{k}$ after a successful inertial step and reset it to a base value otherwise, a common heuristic in accelerated optimization.
+Algorithm 1 Heavy-Ball with Restart
+1:Input: Function $f$ with $L$-Lipschitz gradients, stepsize $\alpha$, initialization $\mathbf{x}_{0}\in\mathbb{R}^{n}$, tolerance $\varepsilon>0$
+2:Set $k=0$, $\beta_{0}=0.5$, $r=\infty$
+3:while $r>\varepsilon$ do
+4:  $\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha\mathbf{\nabla}f(\mathbf{x}_{k})+\beta_{k}\left(\mathbf{x}_{k}-\mathbf{x}_{k-1}\right)$
+5:  $\beta_{k+1}=\tfrac{1}{2}(\beta_{k}+1)$
+6:  if $f(\mathbf{x}_{k+1})>f(\mathbf{x}_{k})-\alpha\left(1-\frac{L\alpha}{2}\right)\|\nabla f(\mathbf{x}_{k})\|_{2}^{2}$ then
+7:   $\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha\mathbf{\nabla}f(\mathbf{x}_{k})$
+8:   $\beta_{k+1}=0.5$
+9:  end if
+10:  $r=\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|/\|\mathbf{x}_{k+1}\|$
+11:  $k\leftarrow k+1$
+12:end while
+13: Output: Approximate solution $\mathbf{x}_{k+1}$
+IV Training on Denoising
+We adopt a bilevel optimization strategy to learn the parameters
+$\boldsymbol{\theta}=\{\mathbf{W},\{\mathbf{Q}_{k},\tau_{k}\}_{k=1}^{K},\boldsymbol{\theta}_{\mathbf{f}},\lambda\},$ (19)
+where $\boldsymbol{\theta}_{\mathbf{f}}$ parametrizes the function $\mathbf{f}$ from Section -C.
+To enforce the constraints $\|\mathbf{W}\|_{2}=1$, $\|\mathbf{Q}_{k}\|_{\infty}\leq 1$, and $\tau_{k}>\|\mathbf{Q}_{k}\|_{2}^{2}$, we optimize over unconstrained auxiliary variables mapped to the feasible set via
+$\mathbf{W}$ $=\frac{\tilde{\mathbf{W}}}{\|\tilde{\mathbf{W}}\|_{2}},\quad\mathbf{Q}_{k}=\frac{\tilde{\mathbf{Q}}_{k}}{\max(1,\|\tilde{\mathbf{Q}}_{k}\|_{\infty})},$ (20)
+[…]
+where the spectral norm $\|\tilde{\mathbf{W}}\|_{2}$ is computed via the Fourier transform [19].
+IV-A Bilevel Formulation
 To make our regularizer more universal, we train it on a range of Gaussian-noise levels and denote by $R_{\sigma}$ the regularizer associated with the noise level $\sigma$.
-To keep the model as simple as possible, we share the filters and most other parameters across noise levels, allowing only $\{\mu_{k}\}_{k=1}^{K}$ to depend on $\sigma$.
-We define a function $\boldsymbol{\mu}\colon\mathbb{R}\to\mathbb{R}^{K}$ that concatenates the $\mu_{k}$, such that $\mu_{k}(\sigma)=\mu(\sigma)_{k}$.
-This function is parameterized as
-$\boldsymbol{\mu}(\sigma)=\frac{1}{100}\left(\sigma\cdot\mathbf{1}_{K}+\frac{1}{20}\mathbf{f}(\bar{\sigma})\right)_{+}+10^{-9}\cdot\mathbf{1}_{K},$ (17)
-where $\mathbf{f}\colon\mathbb{R}\to\mathbb{R}^{K}$ is a three-layer neural network.
-The function $(\cdot)_{+}\coloneqq\max(0,\cdot)$ and the addition of a small term are here to ensure that components of $\boldsymbol{\mu}$ stay positive.
-The term $\bar{\sigma}$ refers to a min-max normalization of $\sigma$ such that it is in the interval [2,2][-2,2] in order to fully benefit from the neural network $\mathbf{f}$ whose knots at initialization are mostly located in that range.
-IV Experiments
-IV-A Training on Denoising
-We denote by $\boldsymbol{\theta}$ the vector of all parameters in our model.
-The paired data used for training are $(\mathbf{x}_{m},\mathbf{y}_{m})_{m=1}^{M}$, where each noisy observation is generated as $\mathbf{y}_{m}=\mathbf{x}_{m}+\sigma_{m}\mathbf{n}_{m}$ with Gaussian noise $\mathbf{n}_{m}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and noise levels $\sigma_{m}\sim\mathcal{U}(0,0.2)$.
-The model is trained to minimize the reconstruction loss
-$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{M}\sum_{m=1}^{M}\frac{1}{\sqrt{\sigma_{m}}}\|\mathbf{x}_{m}^{*}-\mathbf{x}_{m}\|_{1},$ (18)
-where $\mathbf{x}_{m}^{*}\in\operatorname{prox}_{\lambda R_{\sigma_{m}}}(\mathbf{y}_{m})$.
-The weighting by $1/\sqrt{\sigma_{m}}$ ensures that the model gives sufficient importance to low noise levels during training.
-We follow a bilevel strategy to optimize (18).
-The inner optimization corresponds to the solution $\mathbf{x}^{*}_{m}$ to the proximal problem in (2), as found through accelerated gradient descent with restarts, as described in Algorithm 1.
-The inner loop terminates when the relative change of the iterates falls below $10^{-4}$ or after 300 iterations.
-The outer optimization aims to learn $\boldsymbol{\theta}$ by minimizing the empirical loss (18).
-This is done using gradient-based optimization with implicit differentiation to compute the required gradients within the inner-optimization process.
-We use the torchdeq library [24] and apply 25 steps of the Broyden algorithm [25] to approximate the gradient of the fixed-point solution.
+We use a dataset of paired samples $(\mathbf{x}_{m},\mathbf{y}_{m})_{m=1}^{M}$, where $\mathbf{x}_{m}$ is the reference image and $\mathbf{y}_{m}=\mathbf{x}_{m}+\sigma_{m}\mathbf{n}_{m}$ with $\mathbf{n}_{m}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{m}\sim\mathcal{U}(0,0.2)$.
+This yields the bilevel optimization problem
+$\min_{\boldsymbol{\theta}}\quad$ $\mathcal{L}(\boldsymbol{\theta})=\frac{1}{M}\sum_{m=1}^{M}\frac{1}{\sqrt{\sigma_{m}}}\|\mathbf{x}_{m}^{*}(\boldsymbol{\theta})-\mathbf{x}_{m}\|_{1}$ (21)
+s.t.
+$\mathbf{x}_{m}^{*}(\boldsymbol{\theta})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\left(\frac{1}{2}\|\mathbf{x}-\mathbf{y}_{m}\|_{2}^{2}+\lambda R_{\sigma_{m}}(\mathbf{x})\right).$ (22)
+The dependence on $\sigma$ is only enforced through the parameters $\{\mu_{k}\}_{k=1}^{K}$, while the filters and other components remain the same across all noise levels.
+We describe their parameterization in Appendix -C.
+The weighting factor $1/\sqrt{\sigma_{m}}$ ensures that the loss remains sensitive to errors at low noise levels. This is a common strategy used in the training of diffusion models [38].
+IV-B Inner Optimization (Forward Pass)
+The inner problem consists of the minimization of (22).
+Since this objective is not necessarily convex, we find an approximate stationary point $\mathbf{x}_{m}^{*}$ by running the heavy-ball with restart (see Algorithm 1) until the relative change of the iterates falls below $10^{-4}$.
+Since $\mathbf{\nabla}f(\mathbf{x}_{m}^{*})\approx 0$, the point $\mathbf{x}_{m}^{*}$ can be viewed as the fixed point of the gradient-step operator
+$\mathbf{x}_{m}^{*}\approx\mathbf{x}_{m}^{*}-\alpha\mathbf{\nabla}f(\mathbf{x}_{m}^{*})=\mathcal{T}(\mathbf{x}_{m}^{*}).$ (23)
+IV-C Outer Optimization (Backward Pass)
+We need access to the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}$ to minimize the outer objective (21) via gradient descent. Applying the chain rule, we have that
+$\nabla_{\boldsymbol{\theta}}\mathcal{L}=\sum_{m=1}^{M}\left(\frac{\partial\mathbf{x}_{m}^{*}}{\partial\boldsymbol{\theta}}\right)^{T}\nabla_{\mathbf{x}_{m}^{*}}\mathcal{L}.$ (24)
+As shown in the deep equilibrium framework [5], one can compute the gradient of $\mathbf{x}_{m}^{*}$ with respect to $\boldsymbol{\theta}$ using the implicit function theorem applied to (23) as
+$\frac{\partial\mathbf{x}_{m}^{*}}{\partial\boldsymbol{\theta}}=\left(\mathbf{I}-\mathbf{J}_{\mathcal{T}}(\mathbf{x}_{m}^{*})\right)^{-1}\frac{\partial\mathcal{T}}{\partial\boldsymbol{\theta}},$ (25)
+where $\mathbf{J}_{\mathcal{T}}$ is the Jacobian of the fixed-point operator with respect to $\mathbf{x}$.
+The determination of the inverse in (25) is computationally prohibitive.
+Instead, we approximate it using 25 steps of the Broyden algorithm [7], as implemented in the torchdeq library [17].
+This allows us to backpropagate through the equilibrium point without unrolling the entire optimization trajectory, which results in significant reduction of the memory usage.
+IV-D Implementation Details
 The training is conducted on a dataset of 238,400 small patches of size $\left(40\times 40\right)$, extracted from 400 images from the BSD500 dataset [2] and scaled to $[0,1]$.
-This number of training images is relatively small compared to typical deep-learning-based methods, which often rely on significantly bigger datasets to achieve good generalization [27].
-We use the ADAM optimizer with a batch size of 128, setting the learning rate to 0.005 for the filters $\mathbf{W}$ and the function $\boldsymbol{\mu}$, while all other parameters use a learning rate of 0.05.
-The model is trained for 5000 steps, and the learning rates are decayed by a factor of 0.75 every 500 steps.
-The full training procedure takes around 5.5 hours on a Tesla V100 GPU.
-We report the denoising performance of MFoE in Table I, using $K=15$ potentials of input dimension $d=4$ with filters of size $\left(11\times 11\right)$.
-For comparison purposes, we include TV [8], the WCRR [5] and Prox-DRUNet [16].
-Note that the WCRR results differ slightly from those in the original paper, as we trained it on noise levels ranging from 0 to $50/255$, whereas the original implementation was ranging from 0 to $30/255$.
-We achieved this by modifying only the splines $s_{\alpha_{i}}$ from [5, Equation (17)], increasing the number of knots from 11 to 18 to maintain interval lengths as close as possible to the original configuration.
+The number of training images is relatively small compared to typical deep-learning-based methods, which often rely on significantly bigger datasets to achieve good generalization [48].
+We use the ADAM optimizer with a batch size of 128 and set the learning rate to 0.005 for the filters $\mathbf{W}$ and the function $\boldsymbol{\mu}$, while all other parameters use a learning rate of 0.05.
+The model is trained for 5,000 steps and the learning rates are decayed by a factor of 0.75 every 500 steps.
+The full training procedure takes about 5 hours on a Tesla V100 GPU.
+IV-E Results and Ablation Studies
+We report the denoising performance of MFoE in Table I on the grayscale-converted BSD68 [25], McMaster [50] and Set14 [47] datasets.
+We use $K=15$ potentials of input dimension $d=4$ with filters of size $\left(11\times 11\right)$.
+For comparison purposes, we include TV [35], the WCRR [19], and Prox-DRUNet [21].
+The TV regularizer is a classic baseline with a very simple architecture.
+Prox-DRUNet is the most advanced deep-learning-based regularizer.
+It involves three order of magnitudes more parameters than the MFoE.
+Note that the WCRR results differ slightly from those in the original paper because we trained it on noise levels ranging from 0 to $50/255$, whereas the original implementation was ranging from 0 to $30/255$.
+We achieved this by modifying only the splines $s_{\alpha_{i}}$ from [19, Equation (17)], increasing the number of knots from 11 to 18 to maintain interval lengths as close as possible to the original configuration.
 As mentioned in Section III-B, we add the relaxed version of WCRR (WCRR-free) and an alternative version of our model based on the $\ell_{2}$-norm (MFoE-$\ell_{2}$).
-More details about their implementation can be found in Appendices -B and -C.
-Relaxing the 1-weak convexity constraint in WCRR does not yield performance improvements, except at the highest noise level.
-In contrast, the proposed MFoE model consistently outperforms the univariate baselines (WCRR and WCRR-free) as well as the multivariate variant MFoE-$\ell_{2}$.
-Notably, all four models employ the same number of filters $\left(Kd=60\right)$ and have roughly the same number of parameters, which ensures that the observed performance gains are not due to increased model capacity but result instead from the greater expressiveness of the proposed multivariate potentials.
+More details about their implementation can be found in Appendices -D and -E.
+The relaxation of the 1-weak convexity constraint in WCRR yields clear improvements in performance, as evidenced by WCRR-free consistently outperforming the constrained WCRR baseline.
+MFoE-$\ell_{2}$ achieves results comparable to the unconstrained univariate model.
+However, the proposed MFoE model consistently outperforms all univariate baselines (WCRR and WCRR-free) as well as the MFoE-$\ell_{2}$ variant, which confirms the benefits of the learned multivariate $\ell_{\infty}$-based potentials.
+Notably, all four models employ the same number of filters $\left(Kd=60\right)$ and have roughly the same number of parameters, which ensures that the observed performance gains are not due to increased model capacity but result instead from the greater expressivity of the proposed multivariate potentials.
+Although Prox-DRUNet maintains a performance lead, the proposed MFoE reduces this margin while offering a principled, interpretable architecture backed by formal convergence guarantees.
 We present in Table II the performance of the MFoE as we increase the number of filters from $Kd=60$ to 240.
 Unlike the WCRR, which saturates at 60 filters, our model continues to improve and reaches a performance plateau around 180 filters.
-To ensure fair comparisons throughout the paper, we adopt the 60-filter variant as our default model.
-We further investigate the tradeoff between $K$ and $d$ in Table VII, keeping the total number of filters fixed at $Kd=60$.
+To ensure fair comparisons with the WCRR throughout the paper, we adopt the 60-filter variant as our default model.
+We further investigate the tradeoff between $K$ and $d$ in Table VII, keeping the total number of filters fixed at the product $Kd=60$.
 We observe a significant improvement when moving from $d=1$ to $d=2$, with performance peaking at $d=4$.
 Beyond that, performance degrades steadily, with $d=60$ performing worse than even the univariate case $d=1$.
-TABLE I: Denoising performance in terms of noise level on the BSD68 test set.
-$\sigma=15/255$ $\sigma=25/255$ $\sigma=50/255$ Parameters
-TV 29.90 27.48 24.83 1
-WCRR 31.20 28.68 25.68 $1.4\cdot 10^{4}$
-WCRR-free 31.18 28.68 25.78 $1.4\cdot 10^{4}$
-MFoE-$\ell_{2}$ 31.22 28.71 25.80 $1.4\cdot 10^{4}$
-MFoE 31.32 28.84 25.92 $1.4\cdot 10^{4}$
-Prox-DRUNet 31.70 29.18 26.13 $1.7\cdot 10^{7}$
+We attribute the drop in performance to the nature of the $l_{\infty}$-norm.
+The gradient is nonzero only for the filter with the maximal response within each group.
+Fewer filters get updated at each step as $K$ is decreased.
+This dynamic can cause filters that rarely achieve the maximal response to stagnate at their initialization.
+TABLE I: Denoising performance (PSNR / SSIM) on the BSD68, McMaster, and Set14 datasets.
+BSD68 McMaster Set14
+Method $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ Params.
+TV 29.90 / 0.911 27.48 / 0.857 24.83 / 0.771 31.51 / 0.930 29.07 / 0.888 26.28 / 0.822 30.19 / 0.915 27.71 / 0.864 24.86 / 0.782 1
+WCRR 31.20 / 0.933 28.68 / 0.890 25.68 / 0.806 33.24 / 0.952 30.61 / 0.920 27.23 / 0.849 31.74 / 0.938 29.19 / 0.900 25.93 / 0.817 $1.4\cdot 10^{4}$
+WCRR-free 31.18 / 0.932 28.68 / 0.888 25.77 / 0.807 33.32 / 0.953 30.77 / 0.923 27.53 / 0.861 31.76 / 0.937 29.27 / 0.899 26.12 / 0.821 $1.4\cdot 10^{4}$
+MFoE-$\ell_{2}$ 31.21 / 0.933 28.71 / 0.889 25.80 / 0.809 33.32 / 0.953 30.77 / 0.923 27.50 / 0.860 31.77 / 0.938 29.28 / 0.900 26.08 / 0.822 $1.4\cdot 10^{4}$
+MFoE 31.32 / 0.935 28.84 / 0.892 25.93 / 0.812 33.53 / 0.955 31.02 / 0.926 27.73 / 0.865 31.96 / 0.940 29.54 / 0.904 26.35 / 0.827 $1.4\cdot 10^{4}$
+Prox-DRUNet 31.70 / 0.940 29.18 / 0.898 26.14 / 0.815 33.92 / 0.957 31.32 / 0.927 27.97 / 0.860 32.50 / 0.945 30.06 / 0.911 26.83 / 0.835 $1.7\cdot 10^{7}$
 TABLE II: Denoising performance in terms of noise level on the BSD68 test set.
@@
 MFoE ($K=60$, $d=4$) 31.39 28.90 25.97
-IV-B Visualization and Analysis of the Learned Model
-We display the filters of the $(K=15,d=4)$ model in Figure 2.
-Filters within the same group tend to exhibit similar and complementary structures, while filters across different groups often vary significantly.
-Since the potential function depends primarily on the filter that produces the strongest response, this behavior suggests a form of invariance in the model—possibly to orientation, scale, or other transformations.
+V Visualization and Analysis of the Learned Model
+A qualitative comparison of our denoising results with the univariate baselines reveals that our approach performs particularly well on periodic patterns, as shown in Figure 4 on the zebra stripes.
 Figure 2: Learned filters on the denoising task with $K=15$ and $d=4$. The black lines split the filters into the 15 different groups of 4.
+We attribute this improvement to the mechanics of quadrature filtering.
+Consider a pure harmonic signal $x(t)=\cos(\omega_{0}t)$.
+A single linear filter $h$ transforms this signal into $(h*x)(t)=A\cos(\omega_{0}t+\varphi)$, where $A=|H(\omega_{0})|$ and $\varphi=\arg(H(\omega_{0}))$ denote the spectral magnitude and phase, respectively.
+Because the response of the filter oscillates, the regularization penalty induced by a univariate potential will fluctuate spatially.
+This behavior may result in an uneven signal reconstruction.
+To avoid this effect, one may use a pair of quadrature filters $(h_{1},h_{2})$ [1].
+These are built such that $|H_{1}(\omega_{0})|=|H_{2}(\omega_{0})|=A$ and $\arg(H_{2}(\omega_{0}))=\arg(H_{1}(\omega_{0}))+\frac{\pi}{2}$.
+By combining their output via a square root, we get
+$\sqrt{(h_{1}*x)^{2}+(h_{2}*x)^{2}}$
+$\quad=\sqrt{A^{2}\cos^{2}(\omega_{0}t+\varphi)+A^{2}\sin^{2}(\omega_{0}t+\varphi)}$
+$\quad=A.$ (26)
+This is a configuration that cannot be achieved by a univariate architecture.
+Figure 3: A pair of learned filters from the $\left(K=30\text{ and }d=2\right)$ model.
+In the top row, a red frame highlights a location where the two filters differ most markedly.
+The magnitude of the Fourier transforms are shown in the bottom row.
+The filters of the $(K=15,d=4)$ model displayed in Figure 2 exhibit similar and complementary structures within the same group.
+Those filters can work together to extract meaningful patterns in schemes that are reminiscent of quadrature-filter pairs and, more generally, steerable filterbanks [16, 41].
+We show a group of two filters and their Fourier magnitudes from the $(K=30,d=2)$ model in Figure 3.
+Those two filters are almost identical except for a horizontal shift of a group of two pixels, as highlighted by the red frame.
+Because of this, they have very similar Fourier magnitudes, which suggests some quadrature mechanism within this group.
 Figure 4: Denoising results for $\sigma=0.2$, shown on the full image and on a zoomed-in textured region. The reported metrics are (PSNR, SSIM).
-We also plot a learned potential of the $(K=30,d=2)$ model along with its two corresponding filter in Figure 4 and show two additional ones in Section -E.
-Figure 4: Learned potential function and the corresponding filters.
-A direct corollary of Theorem 1 is that the minimizers of our regularizer lie in the null space of $\mathbf{W}$.
+We plot the learned potential of those same two filters in Figure 5.
+Visually, the learned potential resembles a rotated $\ell_{p}$-norm with $p<1$, a geometry characteristic of antisparsity [14].
+The regularizer penalizes correlated responses less severely than independent ones.
+This property favors joint activations and avoids the detrimental periodic nature of individual filter responses via a coupling similar to that of quadrature filters.
+Figure 5: Learned potential function and the corresponding filters. The axes correspond to the responses of the two filters.
+A corollary of Theorem 1 is that the minimizers of our regularizer lie in the null space of $\mathbf{W}$.
 In Figure 6, we show the impulse response of the filter that corresponds to $\mathbf{W}^{T}\mathbf{W}$, along with the magnitude of its Fourier transform.
 While the impulse response is close to a Kronecker impulse, it cannot be exactly so as it must sum up to zero due to the zero-mean constraint on all filters in $\mathbf{W}$.
-The frequency response reveals that only the zero-frequency coefficient vanishes, which suggests that the null space of $\mathbf{W}$ is one-dimensional and consists of constant images.
-However, this is not entirely accurate: because our filters are zero-padded, $\mathbf{W}^{T}\mathbf{W}$ is not perfectly diagonalized by the discrete Fourier transform—although the approximation is very close, as shown in [28].
-It is also intuitive that nonzero constant images are not in the null space of $\mathbf{W}$, as the discontinuities they create at the boundaries result in nonzero filter responses.
-To complete the analysis, we run power iterations on $\left(\mathbf{I}-\mathbf{W}^{T}\mathbf{W}\right)$ in float64 precision to accurately estimate the smallest singular value of $\mathbf{W}$.
-The result, in excess of $2.5\cdot 10^{-3}$, shows that the null space of $\mathbf{W}$ is empty.
-Consequently, our regularizer admits a unique global minimum, the zero-valued image.
-A qualitative comparison of denoising results between our model and univariate baselines reveals that our approach performs particularly well in high-frequency regions that resemble textures as shown in Figure 3.
-This observation aligns with techniques used in classic texture-classification frameworks [29, 30], where features are derived from the strongest response across a bank of filters with varying orientations and scales.
+The frequency response reveals that only the zero-frequency coefficient vanishes.
+This suggests that the null space of $\mathbf{W}$ is one-dimensional and consists of constant images.
 Figure 6: Impulse and frequency response of $\mathbf{W}^{T}\mathbf{W}$.
@@
 The image size on the right is $\left(1500\times 1500\right)$, which accomodates for some amount of zero-padding in the Fourier transform.
-IV-C Inverse Problems
-We benchmark MFoE for image deblurring, CS-MRI, and CT. For this, we simulate measurements as
+VI Inverse Problems
+We benchmark MFoE for image deblurring, CS-MRI, and CT. For each task, we simulate measurements as
 $\mathbf{y}=\mathbf{Hx}+\mathbf{w},$ (27)
-where $\mathbf{H}$ is the linear operator that corresponds to the imaging modality and $\mathbf{w}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathbf{w}}\mathbf{I})$.
-We compare our learned regularizer with TV, the WCRR, and Prox-DRUNet.
-Image reconstruction with these methods requires the minimization of the objective (1).
-This is done via an iterative optimization algorithm which is terminated when the relative difference of the iterates is below $10^{-5}$ or when the number of iterations exceeds 1000.
+where $\mathbf{H}$ is the linear operator corresponding to the imaging modality and $\mathbf{w}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathbf{w}}^{2}\mathbf{I})$.
+We compare our learned regularizer with TV, WCRR, and Prox-DRUNet.
+To reconstruct the images, we minimize the objective (1) using the solvers recommended in the respective original works; namely, proximal gradient descent for TV, safeguarded accelerated gradient descent [19] for WCRR, Douglas-Ratchford splitting [12] for Prox-DRUNet, and the heavy-ball method with restart (Algorithm 1) for MFoE.
+In all cases, the algorithm is terminated when the relative difference between consecutive iterates falls below $10^{-5}$ or when the number of iterations exceeds 1,000.
 To deploy the models, we tune their hyperparameters on a validation set using the coarse-to-fine grid search from [18].
@@
 For the other methods, we tune $\lambda$ and the noise level $\sigma$.
-IV-C1 Deblurring
-We evaluate the frameworks on a deblurring task, where $\mathbf{H}$ is a convolution operator.
-We took from [31] three kernels of size (25 $\times$ 25) shown on the top of Table III.
-One kernel is isotropic Gaussian and the other two are anisotropic kernels that correspond to a motion blur.
-We evaluate the noise levels $\sigma_{\mathbf{w}}\in\{0.01,0.03\}$.
-Each method is initialized with the observed blurred and noisy measurements, the hyperparameters are tuned on the Set12 dataset, and the performances are reported on the BSD68 dataset.
-We report the performances in Table III.
-The MFoE outperforms the WCRR in every instance and manages to match the performances of Prox-DRUNet for the isotropic blur.
+VI-A Deblurring
+We apply the frameworks to a deblurring task, where $\mathbf{H}$ is a convolution operator.
+We take three kernels of size (25 $\times$ 25) from [24], as shown on the top of Table III.
+The first kernel is isotropic Gaussian while the other two are anisotropic and model a motion blur.
+We consider the noise levels $\sigma_{\mathbf{w}}\in\{0.01,0.03\}$.
+Each method is initialized with the observed blurred and noisy measurements.
+The hyperparameters are tuned on the Set12 [49] dataset and the performance is reported in Table III on the BSD68 dataset.
+The MFoE outperforms the WCRR in every instance and manages to match the performance of Prox-DRUNet for the isotropic blur.
 Examples of deblurring are provided in Figure 7.
 There, we can see in the zoomed part that Prox-DRUNet makes a somewhat bolder prediction than the other models.
-This results in a higher SSIM but a slightly lower PSNR than MFoE, despite its blurrier prediction.
-TABLE III: PSNR values for the deblurring on the BSD68 dataset.
+This results in a higher SSIM but a lower PSNR than MFoE, despite the blurrier prediction of MFoE.
+TABLE III: Quantitative results (PSNR / SSIM) for deblurring on the BSD68 dataset.
 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 $\sigma_{\mathbf{w}}=0.01$ $\sigma_{\mathbf{w}}=0.03$ $\sigma_{\mathbf{w}}=0.01$ $\sigma_{\mathbf{w}}=0.03$ $\sigma_{\mathbf{w}}=0.01$ $\sigma_{\mathbf{w}}=0.03$
-TV 26.74 25.62 29.16 26.17 29.36 26.05
-WCRR 27.25 26.09 30.26 26.88 30.38 26.75
-MFoE 27.36 26.18 30.43 27.14 30.65 27.04
-Prox-DRUNet 27.34 26.17 30.55 27.24 30.78 27.15
+TV 26.74 / 0.846 25.62 / 0.803 29.16 / 0.902 26.17 / 0.820 29.36 / 0.904 26.05 / 0.815
+WCRR 27.25 / 0.859 26.09 / 0.819 30.26 / 0.920 26.88 / 0.838 30.38 / 0.922 26.75 / 0.833
+MFoE 27.36 / 0.860 26.18 / 0.819 30.43 / 0.921 27.14 / 0.848 30.65 / 0.926 27.04 / 0.844
+Prox-DRUNet 27.34 / 0.859 26.17 / 0.820 30.55 / 0.924 27.24 / 0.852 30.78 / 0.928 27.15 / 0.848
 Figure 7: Reconstruction for a blur with the third kernel and $\sigma_{\mathbf{w}}=0.03$. The reported metrics are (PSNR, SSIM).
-IV-C2 Compressed-Sensing MRI
-We use fully sampled knee images of size $\left(320\times 320\right)$ from the fastMRI dataset [32] as ground truth, including scans with (PDFS) and without (PD) fat suppression.
-The corresponding measurements are obtained by a subsampling of the 2D Fourier transform ($k$-space) using a Cartesian sampling mask defined by two parameters: the acceleration factor $M_{\text{acc}}$ and the center fraction $M_{\text{cf}}$.
+VI-B Compressed-Sensing MRI
+We use fully sampled real-valued knee images of size $\left(320\times 320\right)$ from a fastMRI dataset [46] as ground truth, including scans with (PDFS) and without (PD) fat suppression.
+The corresponding measurements are obtained by a subsampling of the 2D Fourier transform ($k$-space).
+The Cartesian sampling mask is defined by two parameters: the acceleration factor $M_{\text{acc}}$; and the center fraction $M_{\text{cf}}$.
 Specifically, the central $\lfloor 320M_{\text{cf}}\rfloor$ columns of $k$-space (those corresponding to low frequencies) are fully retained, while the remaining columns are uniformly subsampled, which results in a total of $\lfloor 320/M_{\text{acc}}\rfloor$ selected columns.
-Gaussian noise with standard deviation $\sigma_{\mathbf{w}}=0.01$ is added to the real and the imaginary part of the measurements.
+Gaussian noise with standard deviation $\sigma_{\mathbf{w}}=0.01$ is added to the real and the imaginary parts of the measurements.
 Each method is initialized with the zero-filled image, the hyperparameters are tuned on a validation set of 10 images, and the performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
@@
 Again, the MFoE outperforms the WCRR in every instance.
-It also manages to outperform Prox-DRUNet on the simpler instance with the PD data.
+It also manages to outperform Prox-DRUNet on the $\left(M_{\text{acc}}=4,M_{\text{cd}}=0.08\right)$ instance with the PD data.
 Examples of reconstruction are provided in Figure 8.
-TABLE IV: PSNR values for MRI reconstruction on the fastMRI dataset.
-$\left(M_{\text{acc}}=4,M_{\text{cf}}=0.08\right)$ $\left(M_{\text{acc}}=8,M_{\text{cf}}=0.04\right)$
+TABLE IV: Quantitative results (PSNR / SSIM) for MRI reconstruction on the fastMRI dataset.
+$(M_{\text{acc}}=4,M_{\text{cf}}=0.08)$ $(M_{\text{acc}}=8,M_{\text{cf}}=0.04)$
 PD PDFS PD PDFS
-Zero-fill 27.43 29.62 23.49 27.00
-TV 33.23 32.70 26.97 29.42
-WCRR 35.10 34.18 29.32 31.00
-MFoE 35.40 34.21 30.26 31.20
-Prox-DRUNet 35.32 34.53 30.97 31.66
+Zero-fill 27.43 / 0.843 29.62 / 0.875 23.49 / 0.764 27.00 / 0.833
+TV 33.23 / 0.920 32.70 / 0.905 26.97 / 0.835 29.42 / 0.858
+WCRR 35.10 / 0.950 34.18 / 0.928 29.32 / 0.892 31.00 / 0.888
+MFoE 35.40 / 0.952 34.21 / 0.928 30.22 / 0.903 31.21 / 0.888
+Prox-DRUNet 35.32 / 0.950 34.53 / 0.931 30.97 / 0.904 31.66 / 0.894
 Figure 8: Reconstruction for 8-fold subsampling and PDFS data. The reported metrics are (PSNR, SSIM).
-IV-C3 Computed Tomography
-We evaluate our method on the LoDoPaB-CT dataset [33], which contains ground-truth images derived from 2D slices of reconstructions in the LIDC/IDRI database [34], each of size $\left(362\times 362\right)$.
+VI-C Computed Tomography
+We evaluate our method on the LoDoPaB-CT dataset [23], which contains ground-truth images derived from 2D slices of reconstructions in the LIDC/IDRI database [3], each one of size $\left(362\times 362\right)$.
 The measurement operators are defined by parallel-beam geometry and implemented using the ASTRA Toolbox [42, 29].
@@
 We consider three acquisition settings with 60, 40, and 20 projection angles, respectively.
-Each method is initialized with the filtered-back projected (FBP) image, and the hyperparameters are tuned on a validation set of 10 images.
+Each method is initialized with the filtered-back-projected (FBP) image, and the hyperparameters are tuned on a validation set of 10 images.
 The performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
 The performance is presented in Table V. There too, MFoE outperforms WCRR in all tested configurations. Examples of reconstruction are provided in Figure 9.
-TABLE V: PSNR values for CT reconstruction on the LoDoBaP dataset.
+TABLE V: Quantitative results (PSNR / SSIM) for CT reconstruction on the LoDoBaP dataset.
 60 angles 40 angles 20 angles
-FBP 28.53 25.87 21.41
-TV 33.74 32.58 30.11
-WCRR 35.21 33.87 31.08
-MFoE 35.45 34.22 31.54
-Prox-DRUNet 35.73 34.61 32.12
+FBP 28.53 / 0.796 25.87 / 0.691 21.41 / 0.494
+TV 33.74 / 0.917 32.58 / 0.903 30.11 / 0.867
+WCRR 35.21 / 0.924 33.87 / 0.912 31.08 / 0.880
+MFoE 35.40 / 0.925 34.21 / 0.914 31.48 / 0.885
+Prox-DRUNet 35.73 / 0.927 34.61 / 0.917 32.12 / 0.893
 Figure 9: Reconstruction for CT with 20 angles and 256 line detectors. The reported metrics are (PSNR, SSIM).
-IV-C4 Time Comparisons
+VI-D Statistical Significance
+Across all our inverse problem experiments, we observe that WCRR outperforms TV, MFoE outperforms WCRR, and Prox-DRUNet usually outperforms MFoE.
+We assess the validity of this ranking through Wilcoxon signed-rank tests [44] on the PSNR and SSIM scores across the entire test set, with a strict significance threshold of $p<0.001$.
+In total, we perform 26 separate tests, covering the PSNR and SSIM metrics across all 13 experimental modalities (six deblurring, four CS-MRI, and three CT).
+Our testing framework confirms the robustness of this hierarchy.
+WCRR yields statistically significant improvements over TV in every single configuration.
+The transition from WCRR to MFoE is similarly distinct, with MFoE surpassing WCRR in all PSNR evaluations and achieving significant SSIM gains in 10 out of 13 modalities.
+Finally, Prox-DRUNet proves superior to MFoE in the majority of scenarios; statistical significance is established in 19 of the 26 comparisons, with the few exceptions occurring primarily in deblurring and CS-MRI tasks where both models achieved comparable high-fidelity reconstructions.
+VI-E Time Comparisons
 Table VI contains the average inference time per image, measured in seconds, for each inverse problem.
@@
 These results highlight that the performance gains of Prox-DRUNet come at a computational cost.
-In particular, the CT reconstructions with Prox-DRUNet are significantly slower than with other methods, as they require the repeated evaluation of the proximal operator of the data-fitting term—a step that is especially costly in the case of CT.
-By contrast, MFoE is only moderately more demanding than the WCRR in inference time, between 35% and 40% on average across all experiments due to the computational cost of the projection onto $\mathcal{B}_{\ell_{1}}$.
-TABLE VI: Average time per image for all the reconstruction modalities.
+In particular, the CT reconstructions with Prox-DRUNet require the repeated evaluation of the proximal operator of the data-fitting term—a step that is especially costly in the case of CT.
+This results in Prox-DRUNet being significantly slower than the other methods.
+The MFoE and WCRR methods are on average more than 13 times faster than Prox-DRUNet.
+TABLE VI: Average duration [s] per image for all the reconstruction modalities.
 Deblurring CS-MRI CT
@@
 WCRR 6.45 6.89 13.90
-MFoE 5.39 16.48 16.60
+MFoE 5.90 10.94 10.26
 Prox-DRUNet 39.82 62.52 267.42
-Deep-learning-based methods such as Prox-DRUNet sets a high standard in learned regularization by leveraging a large number of parameters and extensive training data which contribute to their good performance across diverse tasks.
+Although deep-learning-based methods such as Prox-DRUNet set a high standard in learned regularization, they do so at the cost of a large number of parameters and extensive training data.
 We believe that MFoE offers an interesting alternative with significantly fewer parameters, fewer training data, faster inference time, and with a reconstruction quality that is close to Prox-DRUNet.
-Together, these methods offer complementary strengths, providing valuable options suited to varying application needs.
-V Conclusion
+Together, these methods offer complementary strength and provide valuable options suited to a variety of application needs.
+VII Conclusion
 We have proposed the MFoE model.
 It is a principled extension of the classic FoE model and incorporates multivariate potential functions via Moreau envelopes of the $\ell_{\infty}$-norm.
-Our method offers improved adaptability across multiple inverse problems.
-The image quality that it achieves is comparable to that of state-of-the-art learned regularizers.
-It offers a clear structured design that is easy to implement with competitive reconstruction times.
+The image quality that it achieves noticeably narrows the performance gap with state-of-the-art learned regularizers while it offers a clear structured design that is easy to implement with competitive reconstruction times.
 Acknowledgements
-The research leading to these results was supported by the European Research Council (ERC) under European Union’s Horizon 2020 (H2020), Grant Agreement - Project No 101020573 FunLearn.
+The research leading to these results was supported by the European Research Council (ERC) under European Union’s Horizon 2020 (H2020), Grant Agreement - Project No. 101020573 FunLearn.